OpenStudy (anonymous):

Will medal! @CGGURUMANJUNATH Simplify the sum. 6. (6x^3+4x-5)+(-3x^3-7x^2+4) a. -18x^3-7x^2+4x-1 b. 3x^3+7x^2+4x+1 c. 9x^3-3x^3+4x-9 d. 3x^3-7x^2+4x-1 Simplify the sum. 7. (5u^3+5u^2+3)+(2u^3-3u+8) a. 3u^3-3u^2+5u-11 b. 3u^3+5u^2-3u+11 c. 7u^3+2u^2-3u+11 d. 7u^3+5u^2-3u+11

OpenStudy (ilovecake):

Do you have any ideas

OpenStudy (michele_laino):

question 6) hint: I can rewrite your sum as: \[6{x^3} + 4x - 5 - 3{x^3} - 7{x^2} + 4 = ...?\]

OpenStudy (ilovecake):

?

OpenStudy (anonymous):

6. A?

OpenStudy (michele_laino):

please add similar terms each other

OpenStudy (michele_laino):

no, what is 6x^3 - 3x^3=...?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

I just got the answer d

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

question 7) hint: I can rewrite your sum as: \[5{u^3} + 5{u^2} + 3 + 2{u^3} - 3u + 8 = ...?\]

OpenStudy (michele_laino):

again, please sum similar terms each other

OpenStudy (anonymous):

And 7 is d as well

OpenStudy (anonymous):

Right? @Michele_Laino

OpenStudy (michele_laino):

that's right!